OpenStudy (anonymous):

each digit 0,1,2,3,4,5,6,7,8, an 9 is writting on a slip of paper and the slips are placed in a hat. if 3 slips of paper are selected at ramdom find the probl that the 3 numbers are greater then 5

OpenStudy (anonymous):

3/10

OpenStudy (anonymous):

i mean 4/10

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It's \[\frac{4 \choose 3}{10 \choose 3}\]

OpenStudy (anonymous):

there are ten choose 3 ways to choose three numbers out of ten ten choose 3 is 10*9*8/3/2=120 . This is your denominator. There are 4 choose 3 ways to choose the 4 of the ten numbers that are greater than 5. 4 choose 3 is 4. answer is 4/120 = 1/30

OpenStudy (anonymous):

Yeah, that's what I said ;p

OpenStudy (anonymous):

yup. couldn't find n choose k in the template.

OpenStudy (anonymous):

I always write equations manually. It's all just latex =) { n \choose k}

OpenStudy (anonymous):

much easier than using the editor

OpenStudy (anonymous):

thanks for the tip. {x\times y}\[\frac{a}{b}\] just testing

OpenStudy (anonymous):

oops didnt work

OpenStudy (anonymous):

You have to be sure your equation is inside the \[ \ ] bits (no space)

OpenStudy (anonymous):

\[x \times y\]

OpenStudy (anonymous):

\[\frac{a}{b}\]

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

\[x\cdot y, \forall x \in N, f(g(x)) = (f\circ g)(x)\]

OpenStudy (anonymous):

Hrm? you are making fractions?

OpenStudy (anonymous):

trying to typed \[\frac{a}{b}\]

OpenStudy (anonymous):

it's working for me.. (I can see a over b)

OpenStudy (anonymous):

it looks good in preview but doesn't post right

OpenStudy (anonymous):

i see no fraction bar!

OpenStudy (anonymous):

\[x\times y\]

OpenStudy (anonymous):

got one

OpenStudy (anonymous):

\[{n \choose k}\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

Sometimes the fraction bar doesn't show up for you when you post stuff. if you go to a different question and come back it can fix itself.